ROTATIONAL ENERGIES IN VARIOUS TORSIONAL LEVELS OF CH_{2}DOH^{a}

\underline{L. H. COUDERT}, A. EL HILALI, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France; L. MARGULÈS, R. A. MOTIYENKO, Laboratoire PhLAM, UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; and S. KLEE, Physikalisch-Chemisches Institut, Justus-Liebig-Universität Gießen, 35392 Gießen, Germany.

Using an approach accounting for the hindered internal rotation of a monodeuterated methyl group,^{b} an analysis of the torsional spectrum of the monodeuterated species of methanol CH_{2}DOH has been carried out recently and led to the assignment of 76 torsional subbands in its microwave, FIR, and IR spectra. Although this approach also allowed us to account for subband centers, the rotational structure of the torsional subbands is not well understood yet. In this paper, we will deal with the rotational energies of CH_{2}DOH. Analyses of the rotational structure of the available subbands^{b} have been performed using the polynomial-type expansion introduced in the case of the normal species of methanol.^{c} For each subband, FIR or IR transitions and a-type microwave lines, within the lower torsional level, were fitted. The frequencies of the latters were taken from previous investigations^{d} or from new measurements carried out from 50 to 950 GHz with the submillimeterwave solid state spectrometer in Lille. Subbands involving lower levels with v_{t} = 0 and K ≥ 3 could be satisfactorily analyzed. For levels characterized by lower K-values, the expansion fails. In the case of the K = 1, v_{t} = 1 level, the frequencies of a-type microwave transitions involving the lower member of the K-type doublet cannot be well reproduced. For K = 0 levels with v_{t} = 1 and 2, a large number of terms is needed in the expansion. We will try to understand why the rotational energies of these levels cannot be reproduced. The results of the analyses will be compared to those obtained with a global approach based on the rotation-torsion Hamiltonian of the molecule.^{e}

aWork supported by the CNES, the INSU program PCMI, and the ANR-08-BLAN-0225 and ANR-08-BLAN-0054 contracts.

bEl Hilali, Coudert, Konov, and Klee, J. Chem. Phys. 135 (2011) 194309.

cIoli, Moruzzi, Riminucci, Strumia, Moraes, Winnewisser, and Winnewisser, J. Mol. Spec. 171 (1995) 130.

dQuade and Suenram, J. Chem. Phys. 73 (1980) 1127; and Su and Quade, J. Mol. Spec. 134 (1989) 290.

eLauvergnat, Coudert, Klee, and Smirnov, J. Mol. Spec. 256 (2009) 204.